OpenStudy (anonymous):

A ladder is leaning against a house. The distance between the base of the ladder to the house is 3 feet. The length of the ladder is 10 feet. In feet, how far above the ground does the ladder touch the house?

OpenStudy (anonymous):

have to apply Pythagoras theorem mostly 3^2 + x^2=10^2 x=root(100-9)=root(91)=9.53 feet